# On irrationals with Lagrange value exactly 33

### v2

@@ 3.4. Uncountably many elements in $Y_{3}(n)$ for $n\in\mathbb{N}_{+}$
One way of proving that there are uncountably many elements in $X_{3}(n)$ is adding exactly $n$ coefficients equal to $3$ at the beginning of $\omega_{X_{3}(0)}$.
@@
Clearly the cuts at the first $n$ positions (at the letter $3$) are bad cuts, so we only need to prove that there are no bad cuts inside the suffix $\omega_{X_{3}(0)}$. The proof proceeds by induction, which is similar to Lemma 3.6. We only need to check that every odd cut inside the suffix $\omega_{X_{3}(0)}$ is good, the rest part of the induction coincide with the proof of Lemma 3.6.
@@
where we are using the fact that $\omega_{0}\in\{1,2\}$ is smaller than $3$ and the $i$-th position is an odd position. ∎
@@
so there are exactly $2$ bad cuts in the prefix $(\alpha_{1}^{T})^{+}$ or $(\beta_{1}^{T})^{+}$, which are
@@
We prove that the cut at the last $2$ of $(\alpha_{n+1}^{T})^{+}$ or $(\beta_{n+1}^{T})^{+}$ is the only bad cut in the suffix $(\beta_{n}^{T})^{j}\alpha_{n}^{T}$ and this will finish the induction. Notice that the word $(\beta_{n}^{T})^{j}\alpha_{n}^{T}$ can also be written in the form $\beta^{T}$ for some $\beta\in P$, so by Lemma 3.6 we know that all the cuts inside $(\beta_{n}^{T})^{j}\alpha_{n}^{T}$ are good except the cut at the last $2$. By Lemma 3.9 we know that we can always write $\alpha_{n+1}$ as $a\theta b$ and $\beta_{n+1}$ as $a\theta^{\prime}b$, where $\theta,\theta^{\prime}$ are palindrome and $\theta$ is a prefix and suffix of $\theta^{\prime}$. So the cut at the last $2$ is always of the form $\hat{\theta}a|b\hat{\theta}\dots$, where $\hat{\theta}=\theta$ or $\theta^{\prime}$ is palindrome. This cut is a bad cut because of the Proposition 2.11 and the inequality that $\lambda(\hat{\theta}a|b\hat{\theta}\dots)>\lambda(\dots\hat{\theta}a|b\hat{\theta}\dots)=3$.
@@
For $n=1$, there is exactly $1$ bad cut inside the infinite word $a\omega_{X_{3}(0)}$, so $a\omega_{X_{3}(0)}\in X_{3}(1)$. For $n\geq 2$, there are exactly $n$ bad cuts inside the infinite word $(\alpha_{n-1}^{T})^{+}\omega_{X_{3}(0)}$, so $(\alpha_{n-1}^{T})^{+}\omega_{X_{3}(0)}\in X_{3}(n)$.
@@ 4. Related Questions
It is also interesting to study the structure of $\widetilde{\mathcal{M}}$. The first observation is that, different from the Lagrange value $k$, the function $\widetilde{m}$ depends on the initial part of the continued fraction, so maybe the spectrum will not be as “good” as the Lagrange spectrum or Markov spectrum. The first element of $\widetilde{\mathcal{M}}$ is no longer $\sqrt{5}$, in fact, it is $\widetilde{m}([1;1,1,\dots])=\frac{3+\sqrt{5}}{2}$. But similarly with $\mathcal{L}$ and $\mathcal{M}$, is the initial part of $\widetilde{\mathcal{M}}$ also a discrete sequence accumulating precisely at $3$? Does $\widetilde{\mathcal{M}}$ has positive Hausdorff dimension? Does $\widetilde{\mathcal{M}}$ have Hall’s ray?

### v3

@@ 3.4. Uncountably many elements in $Y_{3}(n)$ for $n\in\mathbb{N}_{+}$
One way of proving that there are uncountably many elements in $X_{3}(n)$ is adding exactly $n$ coefficients equal to 33 at the beginning of $\omega_{X_{3}(0)}$.
@@
Clearly the cuts at the first $n$ positions (at the letter 33) are bad cuts, so we only need to prove that there are no bad cuts inside the suffix $\omega_{X_{3}(0)}$. The proof proceeds by induction, which is similar to Lemma 3.6. We only need to check that every odd cut inside the suffix $\omega_{X_{3}(0)}$ is good, the rest part of the induction coincide with the proof of Lemma 3.6.
@@
where we are using the fact that $\omega_{0}\in\{1,2\}$ is smaller than 33 and the $i$-th position is an odd position. ∎
@@
so there are exactly 22 bad cuts in the prefix $(\alpha_{1}^{T})^{+}$ or $(\beta_{1}^{T})^{+}$, which are
@@
We prove that the cut at the last 22 of $(\alpha_{n+1}^{T})^{+}$ or $(\beta_{n+1}^{T})^{+}$ is the only bad cut in the suffix $(\beta_{n}^{T})^{j}\alpha_{n}^{T}$ and this will finish the induction. Notice that the word $(\beta_{n}^{T})^{j}\alpha_{n}^{T}$ can also be written in the form $\beta^{T}$ for some $\beta\in P$, so by Lemma 3.6 we know that all the cuts inside $(\beta_{n}^{T})^{j}\alpha_{n}^{T}$ are good except the cut at the last 22. By Lemma 3.9 we know that we can always write $\alpha_{n+1}$ as $a\theta b$ and $\beta_{n+1}$ as $a\theta^{\prime}b$, where $\theta,\theta^{\prime}$ are palindrome and $\theta$ is a prefix and suffix of $\theta^{\prime}$. So the cut at the last 22 is always of the form $\hat{\theta}a|b\hat{\theta}\dots$, where $\hat{\theta}=\theta$ or $\theta^{\prime}$ is palindrome. This cut is a bad cut because of the Proposition 2.11 and the inequality that $\lambda(\hat{\theta}a|b\hat{\theta}\dots)>\lambda(\dots\hat{\theta}a|b\hat{\theta}\dots)=3$.
@@
For $n=1$, there is exactly 11 bad cut inside the infinite word $a\omega_{X_{3}(0)}$, so $a\omega_{X_{3}(0)}\in X_{3}(1)$. For $n\geq 2$, there are exactly $n$ bad cuts inside the infinite word $(\alpha_{n-1}^{T})^{+}\omega_{X_{3}(0)}$, so $(\alpha_{n-1}^{T})^{+}\omega_{X_{3}(0)}\in X_{3}(n)$.
@@ 4. Related Questions
It is also interesting to study the structure of $\widetilde{\mathcal{M}}$. The first observation is that, different from the Lagrange value $k$, the function $\widetilde{m}$ depends on the initial part of the continued fraction, so maybe the spectrum will not be as “good” as the Lagrange spectrum or Markov spectrum. The first element of $\widetilde{\mathcal{M}}$ is no longer $\sqrt{5}$, in fact, it is $\widetilde{m}([1;1,1,\dots])=\frac{3+\sqrt{5}}{2}$. But similarly with $\mathcal{L}$ and $\mathcal{M}$, is the initial part of $\widetilde{\mathcal{M}}$ also a discrete sequence accumulating precisely at 33? Does $\widetilde{\mathcal{M}}$ has positive Hausdorff dimension? Does $\widetilde{\mathcal{M}}$ have Hall’s ray?